# First-Mover Bias in Gradient Boosting Explanations: Mechanism, Detection, and Resolution

### 22 Mar 2026 v1

@@ 7.3 Limitations
  • Scope. The default pipeline targets XGBoost and TreeSHAP. Extension to other model families (e.g., neural networks with KernelSHAP) would require adapting the diversity selection criteria to non-tree-based importance measures. The fit_from_attributions() interface partially relaxes this constraint: any attribution method producing feature-level vectors can plug into the diversity selection and consensus stages without retraining.
@@
  • Ground truth. On real-world data, we can only evaluate stability, not accuracy. The synthetic accuracy metric presupposes equitable credit distribution (Section 5), partially confounding accuracy with equity.
  • Confidence intervals. With $N_{\text{reps}}=20$, the Wilcoxon test is underpowered for the small effect sizes between DASH and Stochastic Retrain. Larger $N_{\text{reps}}$ would strengthen the analysis.
  • Background dataset size. All experiments use $B=100$ background samples for interventional TreeSHAP. For high-dimensional datasets with strong correlation structure, this may be insufficient to capture the joint distribution faithfully. A sensitivity analysis over $B$ is deferred to the journal version.
  • Random forest baseline. Random forests train trees independently, predicting higher baseline explanation stability without any DASH-like pipeline. Including an RF baseline would provide a direct test of the independence principle on a model family that is independent by construction. This comparison is planned for the journal version.
@@ 7.4 Broader implications
First-mover bias is likely not unique to gradient boosting. Any iterative optimization procedure that makes sequential, greedy feature selections—including some neural network training dynamics—may exhibit analogous path-dependent concentration of feature attributions. The independence principle established here provides a general framework for investigating and resolving such effects. Future work will explore whether the mechanism extends to random forests (where trees are already independent, predicting higher baseline stability), neural networks (where gradient-based optimization creates different but potentially analogous path dependencies), and whether partial orders on feature importance can replace point rankings as a more robust representation of explanation structure. A concrete near-term extension is stable feature interaction estimation: averaging TreeSHAP interaction tensors across the DASH ensemble would provide stable pairwise interaction rankings under multicollinearity, a capability that no single-model workflow can offer.
@@ 8 Conclusion
    DASH and Stochastic Retrain achieve identical stability ($0.977$ at $\rho=0.9$; accuracy $d=+0.05$, equity d=0.21d=-0.21, both n.s.), despite differing in every design choice except model independence. This supports the view that independence between explained models is the operative mechanism.
@@
DASH (Diversified Aggregation of SHAP) operationalizes the independence principle with forced feature restriction, diversity-aware model selection, and two novel diagnostic tools—the Feature Stability Index and Importance-Stability Plot—that detect first-mover bias without ground truth. On real-world datasets, DASH improves stability from $0.32$ to $0.93$ (Breast Cancer), from $0.83$ to $0.96$ (Superconductor), and from $0.97$ to $0.98$ (California Housing).
@@
All code, data generators, hyperparameter configurations, and evaluation metrics are publicly available at https://github.com/DrakeCaraker/dash-shap. The complete benchmark can be reproduced via python run_experiments.py with fixed random seeds ($\text{SEED}=42$). The authoritative interactive notebook notebooks/demo_benchmark_6.ipynb provides a checkpointed walkthrough of all experiments; all tables and figures in this paper are sourced from that notebook. Hardware-dependent timing results (Table 9) were measured on Apple M-series silicon, single node, and are reported for relative comparison only.
@@ Appendix B Extended Results and Reproducibility
Full per-repetition importance vectors, significance test results, and ablation data are available in JSON format from the experimental runner. Running python run_experiments.py reproduces all tables and figures. The authoritative notebook notebooks/demo_benchmark_6.ipynb provides a checkpointed walkthrough of all experiments with intermediate results cached for rapid re-execution.
@@ Appendix C Ablation Studies and Computational Cost
Table 8: Epsilon sensitivity at $\rho=0.9$. Performance is robust across a $3\times$ range.
@@
Stability is robust across population sizes $M$: $M=50$ ($0.9727$) $\to$ $M=100$ ($0.9719$) $\to$ $M=200$ ($0.9722$) $\to$ $M=500$ ($0.9722$). Performance is effectively invariant to population size (within 0.001 across $M\in\{50,100,200,500\}$). We use $M=200$ as a conservative default.
Sensitivity plot of stability vs population size M at two correlation levels
Figure 5: Ablation sensitivity: stability as a function of population size $M$ at two correlation levels ($\rho=0.0$ and $\rho=0.9$). Stability is effectively invariant to population size (within 0.001 across $M\in\{50,100,200,500\}$), indicating that $M=200$ is a conservative default. Additional ablation results for $K$, $\varepsilon$, and $\delta$ are reported in Table 8 and the inline text; full multi-panel ablation figures are planned for the journal version.
@@
Table 9 reports wall-clock time for each method at $\rho=0.9$ (20 repetitions, single-threaded timing). DASH’s cost is dominated by training $M=200$ models and computing $K\leq 30$ TreeSHAP explanations.
Table 9: Computational cost at $\rho=0.9$ (20 repetitions). Wall-clock times are hardware-dependent (Apple M-series, single node) and reported for relative comparison.
@@ Appendix D Pre-Specified Success Criteria
    $K_{\text{eff}}$ increases with $\varepsilon$: monotonic $\to$ PASS ($4.0\to 6.5\to 12.2\to 16.2$)
@@
    Nonlinear DGP: DASH >> SB stability at $\rho=0.9$ $\to$ PASS (DASH =0.8734=0.8734 vs. SB =0.8336=0.8336)
@@
    Superconductor: DASH stability >> SB $\to$ PASS ($0.962$ vs. $0.830$)
@@
    California Housing: DASH stability >> SB $\to$ PASS ($0.982$ vs. $0.967$)
@@
    Breast Cancer: DASH stability >> SB $\to$ PASS ($0.930$ vs. $0.317$)

### 06 Apr 2026 v2

@@ 7.2 Practical recommendations
Model governance considerations.
In regulated settings (e.g., banking model risk management under SR 11-7, or transparency obligations under the EU AI Act), DASH is best understood as an explanation auditing layer rather than a replacement for the production model. The deployed system still trains and serves a single XGBoost model; DASH runs offline during model validation to assess whether that model’s SHAP-based feature rankings are trustworthy. The $M$-model population is an audit instrument— analogous to stress-testing a model under varied specifications—not a deployed ensemble. For model risk documentation, we recommend recording: (1) the production model’s SHAP rankings alongside the DASH consensus rankings, (2) the FSI values and IS Plot quadrant assignments as evidence of explanation reliability or instability, and (3) specific features flagged as Quadrant II members, which should be reported as collinear groups rather than individually ranked. DASH itself need not be separately inventoried as a model; it is a diagnostic tool applied to an existing inventoried model, analogous to backtesting or stress testing.
Expected improvement magnitude.
The Breast Cancer result (+0.549+0.549 over Single Best) reflects extreme collinearity (21 feature pairs with $|r|>0.9$). On datasets with moderate collinearity—California Housing ($|r|\approx 0.7$, 8 features)—the stability improvement is +0.009+0.009 (not significant), and DASH’s primary value is the diagnostic output rather than the stability gain. Practitioners should calibrate expectations accordingly: improvements of +0.01+0.010.050.05 on stability are typical for moderately correlated datasets, with larger gains reserved for settings with heavy collinearity.
@@ 7.3 Limitations
  • Scope of the independence principle. The empirical validation in this paper covers interventional TreeSHAP on XGBoost. The companion impossibility theorem (Caraker et al., 2026b) proves that the underlying instability—and the necessity of ensemble-based resolution—holds for any attribution method on any model class exhibiting the Rashomon property under correlation, providing a theoretical basis for expecting the independence principle to generalize. However, direct empirical validation on other configurations (conditional SHAP, LightGBM, CatBoost, neural networks with KernelSHAP) remains future work. The fit_from_attributions() interface partially bridges this gap: any attribution method producing feature-level vectors can plug into DASH’s aggregation stages without retraining (validated on LIME in Appendix G). The conditional SHAP case warrants particular attention: although conditional approaches account for feature dependencies in the reference distribution (Aas et al., 2021), the companion paper proves that switching to conditional SHAP does not escape the impossibility when features have equal causal effects (Caraker et al., 2026b)—the instability is a property of the Rashomon set, not the attribution method. However, the practical dynamics of instability under conditional SHAP may differ from the interventional setting, and empirical validation remains open.
@@
  • Ground truth. On real-world data, we can only evaluate stability, not accuracy. The synthetic accuracy metric presupposes equitable credit distribution (Section 5), partially confounding accuracy with equity. Specifically, the linear DGP assigns equal true importance to all features within a correlated group by construction, aligning the equity and accuracy metrics by design rather than testing them independently. Appendix E examines the asymmetric case where one feature is causally active and its correlate is a passive proxy, directly testing whether DASH over-equalizes.
  • Background dataset size. All experiments use $B=100$ background samples for interventional TreeSHAP. For high-dimensional datasets with strong correlation structure, this may be insufficient to capture the joint distribution faithfully.
@@ 7.4 Broader implications
First-mover bias is likely not unique to gradient boosting. Any iterative optimization procedure that makes sequential, greedy feature selections—including some neural network training dynamics—may exhibit analogous path-dependent concentration of feature attributions. The independence principle established here provides a general framework for investigating and resolving such effects. Our Random Forest results confirm that tree-level independence already yields high stability (Section 7), validating the principle on a model family that is independent by construction. Future work will explore whether the mechanism extends to neural networks (where gradient-based optimization creates different but potentially analogous path dependencies) and whether partial orders on feature importance can replace point rankings as a more robust representation of explanation structure.
The empirical equivalence between DASH and Stochastic Retrain—and the monotonic degradation of dependent methods with $\rho$—is consistent with a formal impossibility result we establish in companion work (Caraker et al., 2026b): no single-model feature ranking can simultaneously be faithful (reflect the model’s attributions), stable (consistent across equivalent models), and complete (rank all feature pairs) when features are collinear. This trilemma holds for any attribution method applied to any model class exhibiting the Rashomon property under correlation—not only gradient boosting. The present work provides the empirical foundation and constructive resolution; the companion theorem, mechanically verified in the Lean 4 proof assistant, proves that ensemble-based approaches like DASH are not merely effective but mathematically necessary for achieving both stability and equity.
The neural network case warrants particular attention. Attribution instability under feature collinearity is well-documented for NNs: small input perturbations produce large ranking changes (Ghorbani et al., 2019), and independently trained models achieving indistinguishable test loss can produce divergent Shapley rankings (D’Amour et al., 2020)—a direct analog of GBDT first-mover bias operating through initialization-driven symmetry breaking rather than residual sequencing. The Rashomon set framework (Fisher et al., 2019; Semenova et al., 2022) predicts that averaging over independent initializations should cancel this bias by the same argument that motivates DASH: the ensemble mean converges toward the attribution implied by the basin center of mass, not any single sample’s idiosyncratic credit assignment. However, two differences temper this prediction. First, NNs possess a wider Rashomon set under high collinearity (D’Amour et al., 2020), likely requiring larger $K$ for equivalent stability gains. Second, gradient-based attribution methods (Integrated Gradients, GradSHAP) introduce baseline-sensitivity variance absent in TreeSHAP (Alvarez-Melis and Jaakkola, 2018), confounding attribution instability with method instability in any NN experiment. A careful extension of DASH to neural architectures would need to control for both factors—a promising but non-trivial direction. A concrete near-term extension is stable feature interaction estimation: averaging TreeSHAP interaction tensors across the DASH ensemble would provide stable pairwise interaction rankings under multicollinearity, a capability that no single-model workflow can offer.
@@ 8 Conclusion
    DASH and Stochastic Retrain achieve equivalent stability ($0.977$ at $\rho=0.9$; accuracy $d=+0.05$, equity d=0.13d=-0.13, both n.s.; TOST confirmed), despite differing in every design choice except model independence. This supports the view that independence between explained models is the operative mechanism. DASH’s diversity-aware selection achieves this with $K_{\text{eff}}\approx 12$ models (vs. $K=30$ for SR) and provides significantly more equitable attribution spreading within correlated groups ($p<0.001$ vs. Random Selection at every $\rho$ level).
@@
DASH (Diversified Aggregation of SHAP) operationalizes the independence principle at two layers: the core operation—averaging attributions across independent models, proved to be the minimum-variance unbiased estimator (Caraker et al., 2026b)—and the pipeline, which adds forced feature restriction, diversity-aware model selection, and two diagnostic tools (the Feature Stability Index and Importance-Stability Plot) that detect first-mover bias without ground truth. On real-world datasets, DASH improves stability from $0.339$ to $0.925$ (Breast Cancer), from $0.840$ to $0.964$ (Superconductor), and from $0.97$ to $0.98$ (California Housing).
@@
All code, data generators, hyperparameter configurations, and evaluation metrics are publicly available at https://github.com/DrakeCaraker/dash-shap. The complete benchmark can be reproduced via python run_experiments_parallel.py with fixed random seeds ($\text{SEED}=42$). The authoritative interactive notebook notebooks/demo_benchmark_7_parallel.ipynb provides a checkpointed walkthrough of all experiments. The canonical results were produced on AWS SageMaker (ml.g5.16xlarge, 64 vCPU, 248 GB RAM); hardware-dependent timing results (Table 11) were measured on Apple M-series silicon, single node, and are reported for relative comparison only.
@@ Appendix B Extended Results and Reproducibility
Full per-repetition importance vectors, significance test results, and ablation data are available in JSON format from the experimental runner. Running python run_experiments_parallel.py reproduces all tables and figures. The authoritative notebook notebooks/demo_benchmark_7_parallel.ipynb provides a checkpointed walkthrough of all experiments with intermediate results cached for rapid re-execution.
@@ Appendix C Ablation Studies and Computational Cost
Table 10: Epsilon sensitivity at $\rho=0.9$ (50 repetitions). Performance is robust across a $3\times$ range.
@@
Stability is robust across population sizes $M$: $M=50$ ($0.9728$) $\to$ $M=100$ ($0.9759$) $\to$ $M=200$ ($0.9767$) $\to$ $M=500$ ($0.9777$). Performance is effectively invariant to population size (within 0.001 across $M\in\{50,100,200,500\}$). We use $M=200$ as a conservative default.
Four-panel ablation sensitivity plot showing stability vs M, K, epsilon, and delta at three correlation levels
Figure 6: Ablation sensitivity: stability as a function of each DASH hyperparameter at three correlation levels ($\rho\in\{0.0,0.9,0.95\}$). Top left: Population size MM—stability is effectively invariant ($\Delta<0.005$ across $M\in\{50,100,200,500\}$). Top right: Selected models KK—saturates at $K\approx 20$. Bottom left: Filter threshold $\varepsilon$—robust across a $3\times$ range (0.030.030.100.10). Bottom right: Diversity threshold $\delta$—sensitive above $0.05$; the default $\delta=0.05$ is at the elbow.
@@
Table 11 reports wall-clock time for each method at $\rho=0.9$ (50 repetitions, single-threaded timing). DASH’s cost is dominated by training $M=200$ models and computing $K\leq 30$ TreeSHAP explanations.
[…]
@@ Appendix D Pre-Specified Success Criteria
    $K_{\text{eff}}$ increases with $\varepsilon$: monotonic $\to$ PASS ($4.0\to 6.6\to 11.9\to 16.1$)
@@
    Nonlinear DGP: DASH >> SB stability at $\rho=0.9$ $\to$ PASS (DASH =0.887=0.887 vs. SB =0.811=0.811)
@@
    Superconductor: DASH stability >> SB $\to$ PASS ($0.964$ vs. $0.840$)
@@
    California Housing: DASH stability >> SB $\to$ PASS ($0.978$ vs. $0.969$; $p=0.063$ n.s. but positive direction)
@@
    Breast Cancer: DASH stability >> SB $\to$ PASS ($0.925$ vs. $0.376$)
@@
Appendix E Asymmetric Causal DGP Check
Motivation.
The linear DGP used in the main experiments assigns equal true importance to all features within a correlated group, which means that DASH’s equity metric (within-group CV) is evaluated in a regime where equity and accuracy are perfectly aligned by construction. A natural concern is whether DASH over-equalizes when one feature is causally active and its correlate is a passive proxy—that is, does aggregation over independent models wash out a legitimate importance asymmetry?
DGP.
We use a two-feature asymmetric causal DGP:
$y=2f_{0}+\varepsilon,\quad f_{1}=\rho f_{0}+\sqrt{1-\rho^{2}}\,z,\quad\varepsilon,z\sim\mathcal{N}(0,\sigma^{2}),$
where $f_{0}$ is causally active and $f_{1}$ is a passive correlate. Ground-truth importance: $f_{0}=1.0$, $f_{1}=0.0$ (normalized). We sweep $\rho\in\{0.5,0.7,0.9,0.95\}$ with N=5, [^0] N=5{,}000 and $N_{\text{reps}}=50$ repetitions.
Metrics.
For each method we report:
  • Stability: mean pairwise Spearman correlation across repetitions (same as the main experiments).
  • Attribution bias: $|\overline{\hat{\phi}}_{0}-\text{true}(f_{0})|$, the absolute deviation of the mean SHAP attribution for $f_{0}$ from its ground-truth importance (after normalizing attributions to sum to 1).
  • Passive leak: mean attribution assigned to $f_{1}$ (ground truth: 0.0); a proxy for over-equalization.
Results.
Table 12 reports results at $\rho=0.9$ (the primary evaluation condition in the main experiments). DASH correctly preserves the causal asymmetry: all methods attribute substantially more importance to the causal feature $f_{0}$ than the passive correlate $f_{1}$. However, DASH’s passive leak (0.089) is higher than Single Best’s (0.068), reflecting a trade-off inherent to ensemble averaging: some models in the ensemble attribute to $f_{1}$, and their contributions are not zeroed out by averaging. Stochastic Retrain shows a similar pattern (0.074). This trade-off increases with $\rho$ (DASH passive leak rises from 0.046 at $\rho{=}0.5$ to 0.173 at $\rho{=}0.95$), and is the expected cost of the variance reduction that DASH provides.
Table 12: Asymmetric causal DGP at $\rho=0.9$, 50 repetitions. “Passive leak” is mean attribution to the causally inert feature $f_{1}$ (ground truth: 0.0). Lower is better for both bias and passive leak. All methods achieve stability = 1.000 (deterministic DGP).
Method Stability Bias ($f_{0}$) Passive leak ($f_{1}$)
Single Best 1.000 0.068 0.068
Stochastic Retrain 1.000 0.074 0.074
DASH (MaxMin) 1.000 0.089 0.089
Large Single Model 1.000 0.084 0.084
Appendix F Two-Tree Analytical Example
To build intuition for first-mover bias, we analyze the simplest possible case: two-step linear boosting on correlated features. The companion paper [Caraker et al., 2026b] proves the formal result—the attribution ratio is exactly $1/(1-\rho^{2})$ for gradient boosting under the proportionality axiom—and verifies it in Lean 4. Here we provide an accessible derivation that isolates the core mechanism.
Setup.
Let $A,B$ be two features with $\text{Cov}(A,B)=\rho$, $\text{Var}(A)=\text{Var}(B)=1$, and target $y=A+\varepsilon$. Consider linear boosting with learning rate $\eta$: at each step we fit $y$ (or its residual) by regressing on a single feature and updating the model with the fitted value scaled by $\eta$.
Step 1.
Suppose the first step selects feature $A$. The fitted coefficient is $\hat{\beta}_{1}=\text{Cov}(y,A)/\text{Var}(A)=1$, so the model after step 1 is $f_{1}(x)=\eta A$. The residual is $r_{1}=y-\eta A=(1-\eta)A+\varepsilon$.
Step 2.
Now consider the gain from each feature on the residual $r_{1}$:
$\text{Gain}(A)$ $=\bigl[\text{Cov}(r_{1},A)\bigr]^{2}/\text{Var}(A)=(1-\eta)^{2},$ (10)
$\text{Gain}(B)$ $=\bigl[\text{Cov}(r_{1},B)\bigr]^{2}/\text{Var}(B)=\rho^{2}(1-\eta)^{2}.$ (11)
Since $\rho<1$, feature $A$ has strictly higher gain on the residual: $\text{Gain}(A)/\text{Gain}(B)=1/\rho^{2}$. The second step selects $A$ again. After two steps, the model is $f_{2}(x)=\eta(2-\eta)A$: only feature $A$ is used, and all SHAP credit goes to $A$.
From gain bias to concentration.
The gain ratio $1/\rho^{2}\approx 1.23$ at $\rho=0.9$ demonstrates the mechanism: after one step, the residual structure encodes which feature was used, creating a bias toward re-selecting that feature.
This simple linear model is deliberately minimal. Two features and linear gain are sufficient to isolate the residual-bias mechanism, but they do not capture the full concentration dynamics of tree-based boosting. In particular, when the DGP is symmetric ($y=\beta\bar{z}_{g}+\varepsilon$ with $\bar{z}_{g}=(A+B)/2$, the paper’s main setting), linear boosting alternates between $A$ and $B$ at each step—the unused feature always has marginally higher gain—so no net concentration arises in this idealization.
In actual XGBoost, three additional factors break this alternation and enable the concentration observed in Figure 1: (1) threshold-based splits create nonlinear gain functions that do not alternate as cleanly as linear regression gains; (2) multi-depth trees create interaction structure where a feature used at the root influences which features are useful at child nodes; and (3) colsample_bytree restricts the available feature set per tree, so the feature that accumulates slightly more early splits (by chance) builds a residual advantage that compounds over hundreds of trees. The interplay of these stochastic and structural effects is what produces the monotonically growing concentration in Figure 1—the linear gain-bias derived above provides the seed, and tree-specific dynamics amplify it.
Independence resolves the compounding.
Because independently trained models make their first-mover selections independently, the compounding runs in different directions for different models. When their SHAP vectors are averaged:
$\bar{\phi}_{A}\approx\tfrac{1}{K}\sum_{k=1}^{K}\phi_{A}^{(k)},$ (12)
models that concentrated on $A$ and models that concentrated on $B$ cancel each other’s arbitrary choices. With $K$ independent models, the variance of the consensus importance due to first-mover effects decays as $O(1/K)$, explaining the stability plateau at $K\approx 20$ observed in the $K$-sweep experiment.
This example identifies the residual gain-bias seed ($1/\rho^{2}$) and the cancellation principle ($O(1/K)$ variance decay). Figure 1 confirms the empirical predictions: (1) concentration grows monotonically with $T$ for a single sequential model, and (2) independent ensembles remain flat regardless of per-model tree count.
Appendix G Attribution-Agnostic Interface Validation
The fit_from_attributions() interface (contribution 5 in Section 1) decouples DASH’s aggregation stages (filtering, diversity selection, consensus, diagnostics) from the attribution method. We validate this claim by applying DASH to LIME [Ribeiro et al., 2016] attributions on the Breast Cancer dataset.
Procedure.
We train $M=30$ XGBoost classifiers with randomly sampled hyperparameters (same search space as the main experiments, colsample_bytree $\in\{0.2,0.3,0.4,0.5\}$). For each model, we compute LIME attributions for $N^{\prime}=30$ observations from the held-out explain set, producing an $(M,N^{\prime},P)$ attribution tensor. We then call pipe.fit_from_attributions(attributions, val_scores) with $K=10$ and relative $\varepsilon=0.05$.
Results.
DASH successfully executes all four post-population stages on the LIME tensor. The consensus top features—worst concave points (0.104), worst texture (0.102), worst area (0.086), worst perimeter (0.073)—draw from the same clinically relevant feature families (concavity, area, perimeter) as the TreeSHAP-based rankings in Table 8, confirming that the diagnostic framework generalizes across attribution methods. MaxMin diversity selection yields $K_{\text{eff}}=3$ (low because all 30 models achieve identical validation accuracy on this dataset, limiting diversity in the filtered pool).
This demonstrates that DASH’s stages 2–5 can operate on any $(M,N^{\prime},P)$ attribution tensor, regardless of whether it was produced by TreeSHAP, LIME, Integrated Gradients, or any other feature-level attribution method. The independence principle—that averaging over independently trained models cancels arbitrary attribution choices—applies to any attribution method that is sensitive to model specification, not only TreeSHAP.